# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ 2.2.3 Weight Alignment
The linear mode connectivity (LMC) property of deep neural networks demonstrates that there is a connected path between multiple local minima of deep neural networks along which the loss remains nearly constant [51, 41, 183, 42, 43]. Numerous studies [47, 131, 42] have shown that two independent models, starting from the same pre-trained model and fine-tuned with different hyper-parameter configurations, typically satisfy LMC. Further, Adilova et al. [3] and Zhou et al. [241] extended the study of LMC to the layer level. The LMC property implies that multiple local minima may be equivalent in the weight space, and different weight configurations of the same model may represent the same functionality. Inspired by this, many works proposed to permute the weights of one model (i.e., $\small\Theta^{(1)}\rightarrow\Pi(\Theta^{(1)})$) to align with the other model $\Theta^{(2)}$ when merging/interpolating two separate models, as illustrated in Figure 3 (b). $\Pi(\cdot)$ denotes a permutation function, and researchers have dedicated efforts to studying effective and efficient permutation strategies for model alignment.
OTFusion [167] and Imfeld et al. [72] adopt optimal transport to soft-align neurons across models. NeuronAlignment [183] introduces an inexpensive heuristic algorithm to approximate the optimal neuron alignment. CCAMerge [64] permutes by maximizing the correlation between linear combinations of neurons. Notably, Git re-basin [5] proposes three methods –activation matching, weight matching, and straight-through estimation– to align (or permute) the weights of models trained on different tasks. Based on the Git re-basin, Peña et al. [140] further incorporate a Sinkhorn-based projection to improve these alignment methods. In addition, MuDSC [215] proposes simultaneously performing model alignment in weight and activation spaces. Unlike heuristic alignment strategies, Deep-Align [133] proposes a learning-based approach to weight alignment, employing a novel learnable architecture that takes two sets of weights as input and outputs a permutation matrix for alignment.
Despite the significant improvement of these alignment algorithms, Jordan et al. [88] argue that the success of these methods depends on the use of normalization layers (e.g., BatchNorm, LayerNorm, etc.) in the model; without these, the performance of the matching algorithms is greatly reduced. The authors call this the “variance collapse” problem and propose the REPAIR method to solve it. Additionally, Crisostomi et al. [29] noted that previous pairwise permutations do not guarantee cycle consistency, making the alignment fragile. They further proposed to globally optimize the permutations of all layers simultaneously at each step. Overall, aligned models experience much less interference or conflict during merging compared to directly merging unaligned models.
2.3 During Merging Methods
@@
In this section, we provide a detailed discussion on how to merge a set of well-trained models. The existing methods can be roughly divided into five categories: basic merging methods (§2.3.1), weighted-based merging methods (§2.3.2), subspace-based merging methods (§2.3.3), routing-based merging methods (§2.3.4), and post-calibration based methods (§2.3.5).
2.3.1 Basic Merging Methods
@@
One of the most straightforward approaches to model merging is to directly weighted average the parameters of multiple models [190, 164], i.e., $\small\Theta^{(merge)}=\sum_{t=1}^{T}\frac{1}{T}\Theta^{(t)}$. However, the performance of simple weight averaging is generally unsatisfactory. Recently, Task Arithmetic [71] introduced the concept of “task vector” (in Figure 4(a)), which represents the model parameter $\Theta^{(t)}$ fine-tuned on task $t$ subtract the pre-trained model parameter $\Theta^{(0)}$, i.e., $\tau_{t}=\Theta^{(t)}-\Theta^{(0)}$. In other words, task vectors are thought to steer the behavior of a neural network meaningfully. For example, multitask learning (MTL) can be accomplished by adding task vectors, forgetting can be achieved by subtracting task vectors, and task analogies can be performed using analogous task vectors. Specifically, when we want the pretrained model to perform MTL, we can add multiple task vectors $\{\tau_{1},\ldots,\tau_{T}\}$ to the pretrained model, i.e., $\small\Theta^{(merge)}=\Theta^{(0)}+\lambda\cdot\sum_{t=1}^{T}\tau_{t}$ in Figure 4(b), where $\lambda$ is a hyperparameter. Conversely, when we want the pretrained model to forget a function $t$, we can subtract the corresponding task vector from pretrained model as Figure 4(c), i.e., $\small\Theta^{(merge)}=\Theta^{(0)}-\tau_{t}$. As shown in Figure 4(d), we can also implement task analogies by task vector analogies, thus enabling zero-shot learning of new tasks. Similarly, PEMs [234] combines Adapters with different capabilities by extending task arithmetic [71] to parameter-efficient fine-tuning settings. However, the performance of basic merging methods is not satisfactory most of the time, especially when the tasks interfere with each other.
@@
Figure 4: An illustration of Task Arithmetic [71]. (a) Definition of the “task vector”, which is the difference between the fine-tuned model and the pre-trained model. (b) Multi-task learning is performed by merging multiple task vectors. (c) Knowledge forgetting is achieved by subtracting the task vector. (d) Analogical task vectors are used to implement task analogies.
2.3.2 Weighted-based Merging Methods
@@
As we all know, different models (or task vectors) represent different functions, and intuitively, different functions have varying degrees of importance. Therefore, advanced weighted-based model merging methods design various clever rules to determine the merging coefficients, as shown in Figure 5(a). For instance, when merging two models $\Theta^{(1)}$ and $\Theta^{(2)}$ (or task vectors $\tau_{1}$ and $\tau_{2}$), the goal of the weighted merging method is to find the optimal coefficients $\lambda_{1}^{*}$ and $\lambda_{2}^{*}$ so that the merged model $\small\Theta^{(merge)}=\lambda_{1}^{*}\Theta^{(1)}+\lambda_{2}^{*}\Theta^{(2)}$ (or $\small\Theta^{(merge)}=\Theta^{(0)}+\lambda_{1}^{*}\tau_{1}+\lambda_{2}^{*}\tau_{2}$) can retain the capabilities of the independent models as much as possible. However, when the number of models is large, it is impractical to use brute-force grid search to find the optimal merging coefficient because of the expensive search cost involved.
@@
To determine the merging coefficient more effectively, Evolutionary-model-merge [6] and Checkpoint Merging [113] efficient searches for the merging coefficients using evolutionary algorithms and Bayesian optimization, respectively. AdaMerging [220] uses gradient descent optimization to learn the merging coefficients by minimizing entropy as a surrogate loss in unlabeled test data. MetaGPT [240] casts the model merging problem as an MTL formalism, where the goal is to minimize the average loss of the merged model and the independent model. It employs local linearization of the model and the orthogonality of the task vectors to derive the optimal merging coefficient $\lambda_{t}^{*}$ for each model $\tau_{t}$ as follows: $\small\lambda_{t}^{*}={\left\|\tau_{t}\right\|^{2}}/{\sum_{k=1}^{T}\left\|\tau_{k}\right\|^{2}}$. SLERP [53] performs spherical interpolation of the parameters of the two models. The interpolated coefficients of $\tau_{1}$ and $\tau_{2}$ are given by $\small\lambda_{1}^{*}=\frac{\sin\left(\left(1-\lambda\right)\cdot\rho\right)}{\sin(\rho)}$ and $\small\frac{\sin(\lambda\cdot\rho)}{\sin(\rho)}$, respectively, where $\small\rho=\arccos\frac{\tau_{1}\cdot\tau_{2}}{\left|\tau_{1}\right|\cdot\left|\tau_{2}\right|}$ denotes the angle between the two task vectors, and $\lambda$ represents the merging coefficient of the initial setting.
@@
The above-sophisticated weighting methods operate at the model (or task) level. It is well known that each layer and even each neuron in a deep neural network model play a significantly different role, and some research has developed more fine-grained weighted merging strategies. For example, Layer-wise AdaMerging [220] and aTLAS [231] adaptively learn different sets of merging coefficients for each layer or module of the model, respectively. RegMean [86] indicates that closed-form solutions (relying on the data statistics provided by the training set) exist for linear layers in model merging, while nonlinear layers can simply perform weight averaging. Other works utilize the Fisher information matrix [44] to assess the importance of parameters when merging [126, 80]. Fisher-Merging [126] performs model merging based on the importance of the parameters in each independent model, that is, $\small\Theta^{(\text{merge})}=\sum_{t=1}^{T}F^{(t)}\Theta^{(t)}/\sum_{t=1}^{T}F^{(t)}$, where $F^{(t)}$ is the diagonal of the Fisher information matrix with respect to task $t$. Fisher-nodes-merging [186] also combines a set of Transformer [191] models based on the Fisher information matrix. MaTS [175] developed a block diagonal approximation for Fisher merging. Daheim et al. [31] linked the inaccuracy of weighted average with gradient mismatch, and further proposed an uncertainty-based algorithm to reduce the matching error, ultimately merging the models based on a second-order Hessian estimation.
@@
Figure 5: (a) An illustration of weighted-based model merging methods. (b) An illustration of the subspace-based merging method, where empty means zero value. (c) An illustration of the routing-based merge method dynamically performs the model merge based on the input.
2.3.3 Subspace-based Merging Methods
Another class of advanced methods transforms models into sparse subspaces for merging, thereby mitigating task interference. The over-parameterized nature of neural networks and the success of model pruning [59, 24] show that removing most of the parameters from the model barely affects its accuracy [216]. This insight opens up new opportunities for model merging, allowing us to remove insignificant neurons from a single model and merge multiple sparse models within the parameter subspace [33, 60], as shown in Figure 5 (b).
@@ 2.3.3 Subspace-based Merging Methods
TIES-Merging [216] proposes to trim each individual model based on parameter magnitudes, retaining only the top 20% of parameters with the highest magnitudes. It further suggests eliminating parameter sign conflicts to reduce interference, and finally merging sparse models using Task Arithmetic [71]. Similarly, Drop And REscale (DARE) [226] also sparsifies by parameter magnitude, and highlights the importance of further performing rescaling on sparse models. In addition to removing the tail parameters with the smallest weight, the Model Breadcrumbs [32] highlight the importance of removing the parameters (outliers) with the largest weights to further reduce noise in model merging and enhance generalization to hyperparameters. TALL-masks [200] creates a mask matrix specific to each task based on a predefined threshold related to independent models. Further, Model Tailor [243] masks unimportant parameters based on the sensitivity of fine-tuned parameters to loss changes and the significance of changes compared to pre-trained parameters. APL [96] proposes estimating parameter importance using causal interventions. Unlike the standard practice of obtaining a single model through model merging, EMR-Merging [68] proposes maintaining a shared model among multiple tasks alongside a sparse task-specific model. In this approach, the value of the shared model at each index is the largest parameter value among all models. In contrast to the mask construction rules of the aforementioned heuristics, Concrete [176] frames mask construction and model merging as a learnable bi-level optimization problem. The outer-level optimizes the mask matrix, while the inner-level merges the model based on the mask matrix and optimizes it using the unlabeled test samples.
2.3.4 Routing-based Merging Methods
The basic, weighted-based, or subspace-based merging methods discussed in §2.3.1, §2.3.2 and §2.3.3 are static merging methods. This means that the merged model remains the same for all samples or tasks. Given that there are differences between input samples/tasks, the model’s ability may vary when processing different samples/tasks. As shown in Figure 5 (c), some works propose to dynamically merge models (or subsets of layers) based on the samples/tasks [105, 130, 179, 121, 91] during the inference phase.
@@ 2.3.4 Routing-based Merging Methods
For a given input, SMEAR [130] first computes a weighted average of the parameters of each expert by using the distribution of router inputs to the expert modules. The advantage of this approach is that it has a similar computational cost to that of a single expert. Twin-Merging [121] also adaptively combines task-shared and task-private knowledge based on routing during the inference phase. Similarly, Weight-Ensembling MoE [179] proposes a dynamic merging Transformer architecture. Specifically, they observed that the parameters of the linear layer in the fine-tuned model changed more dramatically than those of the nonlinear layer, which also significantly impacted the merging performance. Therefore, they use a standard weighted average to merge all modules except the linear layer. The linear layer is dynamically weighted and merged according to the routing network (sample features as the input of the routing, and merging coefficients as the output) during inference. PWE MoE [178] further extends Weight-Ensembling MoE to a multi-objective optimization setting and uses the preference vector as input for routing.
2.3.5 Post-calibration based Methods
@@
Recently, Yang et al. [219] introduce a post-merging method to calibrate merged models. They observed that merged models (across multiple mainstream model merging methods) suffer from representation bias, meaning the representations extracted by the independent and merged models are very different, leading to performance degradation in the merged model. To alleviate this problem, they propose a module called ‘representation surgery’ to calibrate the representation bias. The core idea is to align the representation of the merged model after representation surgery with that of the independent model.
2.4 Theories and Analysis of Model Merging

### v5

@@ 2.2.3. Weight Re-basin or Alignment
The linear mode connectivity (LMC) property of deep neural networks demonstrates that there is a connected path between multiple local minima of deep neural networks along which the loss remains nearly constant (Garipov et al., 2018; Draxler et al., 2018; Tatro et al., 2020; Entezari et al., 2022; Ferbach et al., 2024; Zhao et al., 2025). Numerous studies (Frankle et al., 2020; Nagarajan and Kolter, 2019; Entezari et al., 2022; Zhang et al., 2025; Xu et al., 2024b) have shown that two independent models, starting from the same pre-trained model and fine-tuned with different hyper-parameter configurations, typically satisfy LMC. Further, Adilova et al. (2024) and Zhou et al. (2023) extended the study of LMC to the layer level. The LMC property implies that multiple local minima may be equivalent in the weight space, and different weight configurations of the same model may represent the same functionality. Inspired by this, many works proposed to permute the weights of one model (i.e., $\small\Theta^{(1)}\rightarrow\Pi(\Theta^{(1)})$) to align with the other model $\Theta^{(2)}$ when merging/interpolating two separate models, as illustrated in Figure 3 (c). $\Pi(\cdot)$ denotes a permutation function, and researchers have dedicated efforts to studying effective and efficient permutation strategies for model alignment. OTFusion (Singh and Jaggi, 2020) and Imfeld et al. (2024) adopt optimal transport to soft-align neurons across models. NeuronAlignment (Tatro et al., 2020) introduces an inexpensive heuristic algorithm to approximate the optimal neuron alignment. CCAMerge (Horoi et al., 2024) permutes by maximizing the correlation between linear combinations of neurons. Notably, Git re-basin (Ainsworth et al., 2023) proposes three methods –activation matching, weight matching, and straight-through estimation– to align (or permute) the weights of models trained on different tasks. Based on the Git re-basin, Peña et al. (2023) further incorporate a Sinkhorn-based projection to improve these alignment methods. Unlike heuristic alignment strategies, Deep-Align (Navon et al., 2024) proposes a learning-based approach to weight alignment, employing a novel learnable architecture that takes two sets of weights as input and outputs a permutation matrix for alignment. Despite the significant improvement of these alignment algorithms, Jordan et al. (2023) argue that the success of these methods depends on the use of normalization layers (BatchNorm, LayerNorm, etc.) in the model; without these, the performance of the matching algorithms is greatly reduced. Additionally, Crisostomi et al. (2024) noted that previous pairwise permutations do not guarantee cycle consistency, making the alignment fragile. They further proposed to globally optimize the permutations of all layers simultaneously at each step. Overall, aligned models experience much less interference after merging compared to directly merging unaligned models.
2.3. During Merging Methods
@@
In this section, we provide a detailed discussion on how to merge a set of well-trained expert models. The existing methods can be roughly divided into six categories: basic methods (§2.3.1), weighted-based methods (§2.3.2), subspace-based methods (§2.3.3), optimization-based methods (§2.3.4), routing-based methods (§2.3.5), and other merging methods (§2.3.6).
2.3.1. Basic Merging Methods
@@
One of the most straightforward approaches to model merging is to directly weighted average the parameters of multiple models (Utans, 1996; Shoemake, 1985; Wortsman et al., 2022a), i.e., $\small\Theta^{(merge)}=\frac{1}{T}\sum_{t=1}^{T}\Theta^{(t)}$. Recently, Task Arithmetic (Ilharco et al., 2023) introduced the concept of “task vector” (in Figure 4(a)), which represents the model parameter $\Theta^{(t)}$ fine-tuned on task $t$ subtract the pre-trained model parameter $\Theta^{(0)}$, i.e., $\tau_{t}=\Theta^{(t)}-\Theta^{(0)}$. In other words, task vectors are thought to steer the behavior of a neural network meaningfully. For example, multitask learning (MTL) can be accomplished by adding task vectors, forgetting can be achieved by subtracting task vectors, and task analogies can be performed using analogous task vectors. Specifically, when we want the pretrained model to perform MTL, we can add multiple task vectors $\{\tau_{1},\ldots,\tau_{T}\}$ to the pretrained model, i.e., $\small\Theta^{(merge)}=\Theta^{(0)}+\lambda\cdot\sum_{t=1}^{T}\tau_{t}$ in Figure 4(b), where $\lambda$ is a hyperparameter. Conversely, when we want the pretrained model to forget a function $t$, we can subtract the corresponding task vector from pretrained model as Figure 4(c), i.e., $\small\Theta^{(merge)}=\Theta^{(0)}-\tau_{t}$. As shown in Figure 4(d), we can also implement task analogies by task vector analogies, thus enabling zero-shot learning of new tasks. Similarly, PEMs (Zhang et al., 2023a) combines Adapters with different capabilities by extending task arithmetic to parameter-efficient fine-tuning settings. However, the performance of basic merging methods is not satisfactory most of the time, especially when the tasks interfere with each other.
@@
Figure 4. An illustration of Task Arithmetic (Ilharco et al., 2023). (a) Definition of the “task vector”, which is the difference between the fine-tuned model and the pre-trained model. (b) Multi-task learning is performed by merging multiple task vectors. (c) Knowledge forgetting is achieved by subtracting the task vector. (d) Analogical task vectors are used to implement task analogies.
2.3.2. Weighted-based Merging Methods
@@
Different models (or task vectors) represent different functions, and intuitively, different functions have varying degrees of importance. Therefore, weighted-based model merging methods design various clever rules to determine the merging coefficients (Minut et al., 2025; Dai et al., 2025), as shown in Figure 5(a). For instance, when merging two models $\Theta^{(1)}$ and $\Theta^{(2)}$ (or task vectors $\tau_{1}$ and $\tau_{2}$), the goal of the weighted merging method is to find the optimal coefficients $\lambda_{1}^{*}$ and $\lambda_{2}^{*}$ so that the merged model $\small\Theta^{(merge)}=\lambda_{1}^{*}\Theta^{(1)}+\lambda_{2}^{*}\Theta^{(2)}$ (or $\small\Theta^{(merge)}=\Theta^{(0)}+\lambda_{1}^{*}\tau_{1}+\lambda_{2}^{*}\tau_{2}$) can retain the capabilities of the independent models as much as possible. A brute-force grid search to find the optimal merging coefficients is impractical when the number of models is large.
@@
In order to determine the merging coefficient more effectively, Evolutionary-model-merge (Akiba et al., 2025) and Checkpoint Merging (Liu et al., 2024d) efficient searches for the merging coefficients using evolutionary algorithms and Bayesian optimization, respectively. AdaMerging (Yang et al., 2024c) uses gradient descent optimization to learn the merging coefficients by minimizing entropy as a surrogate loss in unlabeled test data. MetaGPT (Zhou et al., 2024b) casts the model merging problem as an MTL formalism, and it employs local linearization of the model and the orthogonality of the task vectors to derive the optimal merging coefficient $\lambda_{t}^{*}$ for each model $\tau_{t}$ as follows: $\small\lambda_{t}^{*}={\left\|\tau_{t}\right\|^{2}}/{\sum_{k=1}^{T}\left\|\tau_{k}\right\|^{2}}$. STF (Qiu et al., 2025) identifies task-specific features via singular value decomposition (SVD) and formulates the merging process as a linear system to solve for the optimal task-specific merging coefficients. SLERP (Goddard et al., 2024b) performs spherical interpolation of the parameters of the two models. The interpolated coefficients of $\tau_{1}$ and $\tau_{2}$ are given by $\small\lambda_{1}^{*}=\frac{\sin\left(\left(1-\lambda\right)\cdot\rho\right)}{\sin(\rho)}$ and $\small\lambda_{2}^{*}\frac{\sin(\lambda\cdot\rho)}{\sin(\rho)}$, respectively, where $\small\rho=\arccos\frac{\tau_{1}\cdot\tau_{2}}{\left|\tau_{1}\right|\cdot\left|\tau_{2}\right|}$ denotes the angle between the two task vectors, and $\lambda$ represents the merging coefficient of the initial setting. However, SLERP can only merge two models.
@@
The above weighting methods operate at the model (or task) level. It is well known that each layer and even each neuron in a deep neural network model plays a significantly different role, and some research has developed more fine-grained weighted merging strategies (Wang et al., 2025a). For example, Layer-wise AdaMerging (Yang et al., 2024c) and aTLAS (Zhang et al., 2024a) adaptively learn different sets of merging coefficients for each layer or module of the model, respectively. RegMean (Jin et al., 2023) indicates that closed-form solutions (relying on the data statistics provided by the training/validation set) exist for linear layers in model merging, while nonlinear layers can simply perform weight averaging. Other works utilize the Fisher information matrix (Fisher, 1922) to assess the importance of parameters when merging (Matena and Raffel, 2022; Jhunjhunwala et al., 2024a; Thennal et al., 2024; Tam et al., 2023). Fisher-Merging (Matena and Raffel, 2022) performs model merging based on the importance of the parameters in each independent model, that is, $\small\Theta^{(\text{merge})}=\sum_{t=1}^{T}F^{(t)}\Theta^{(t)}/\sum_{t=1}^{T}F^{(t)}$, where $F^{(t)}$ is the diagonal of the Fisher information matrix with respect to task $t$. Daheim et al. (2024) linked the inaccuracy of weighted average with gradient mismatch, and further proposed an uncertainty-based algorithm to reduce the matching error, ultimately merging the models based on a second-order Hessian estimation.
@@
Figure 5. An illustration of (a) weight-based model merging; (b) subspace-based merging, where empty entries denote zeros; (c) optimization-based merging; and (d) routing-based merging that dynamically merges models based on the input.
2.3.3. Subspace-based Merging Methods
@@
Another class of advanced methods transforms models into sparse or low-rank subspaces for merging, thereby mitigating task interference, as shown in Figure 5(b). (i) Sparse Subspace Merging: TIES-Merging (Yadav et al., 2023) proposes to trim each individual model based on parameter magnitudes, retaining only the top 20% of parameters with the highest magnitudes. It further suggests eliminating parameter sign conflicts to reduce interference, and finally merging sparse models using Task Arithmetic (Ilharco et al., 2023). Similarly, DARE (Yu et al., 2024b) also sparsifies by parameter magnitude, and highlights the importance of further performing rescaling on sparse models. In addition to removing the tail parameters with the smallest weight, the Model Breadcrumbs (Davari and Belilovsky, 2023) highlight the importance of removing the parameters (outliers) with the largest weights to further reduce noise in model merging and enhance generalization to hyperparameters. TALL-masks (Wang et al., 2024a) creates a mask matrix specific to each task based on a predefined threshold related to independent models. Further, Model Tailor (Zhu et al., 2024) masks unimportant parameters based on the sensitivity of fine-tuned parameters to loss changes and the significance of changes compared to pre-trained parameters. APL (Kong et al., 2024) proposes estimating parameter importance using causal interventions. PCB-Merging (Du et al., 2024) measures parameter importance within each task as well as the similarity of parameter importance across tasks, and discards parameters with low importance. CALM (Yan et al., 2025) integrates local information that is consistent with the global task consensus by aligning local binary masks with the global consensus and performing conflict-free merging. In contrast to the mask construction rules of the aforementioned heuristics, Concrete (Tang et al., 2023) frames mask construction and model merging as a learnable bi-level optimization problem. The outer-level optimizes the mask matrix, while the inner-level merges the model based on the mask matrix and optimizes it using the unlabeled test samples.
(ii) Low-Rank Subspace Merging: Low-rank subspace methods extract the most informative components via matrix factorization for merging, while discarding less important components that are more likely to cause conflicts. TSV-M (Gargiulo et al., 2025) applies SVD to task vectors, concatenates the left and right singular vectors, and then orthogonalizes them to remove redundant bases. ISO-C (Marczak et al., 2025) further flattens the singular value spectrum of the task matrices. HO-GSVD (Skorobogat et al., 2025) observes a rank-collapse phenomenon in the task-vector space during merging. To address this, it applies a subspace boosting procedure after generalized SVD to preserve the rank of the task vectors. KnOTS (Stoica et al., 2024b) first concatenates all task vectors and performs SVD on the resulting high-dimensional matrix while enforcing a shared left singular matrix across tasks. It then merges the task-specific right singular matrices for different tasks. OPCM (Tang et al., 2025b) projects the new task vector onto the orthogonal subspace of the previous task vectors to enable continual model merging.
2.3.4. Optimization-based Merging Methods.
Optimization-based methods exploit the intrinsic properties of task vectors to formulate an optimization objective, thereby enabling data-free minimization of inter-task interference or information loss after merging, as shown in Figure 5(c). For example, AWD (Xiong et al., 2024) theoretically shows that interference between tasks can be minimized when task vectors are mutually orthogonal. Building on this insight, it proposes an adaptive weight disentanglement scheme to optimize the original task vectors, making them more orthogonal. DOGE (Wei et al., 2025c) formulates model merging as a constrained optimization problem and interprets task vectors as accumulated gradients. It then solves this problem via an adaptive projected gradient descent algorithm, enabling data-free optimization of task vectors to alleviate conflicts. WUDI-Merging (Cheng et al., 2025) shows that the task vectors of a linear layer span an approximate linear subspace of its corresponding inputs, and leverages this property to construct a data-free loss function that minimizes interference among task vectors. DOP (Yang et al., 2025a) combines data-space approximation with orthogonal projection, and employs a data-free loss function to balance stability and plasticity, thereby enabling continual model merging.
2.3.5. Routing-based Merging Methods
@@
The basic, weighted, subspace or optimization-based merging methods discussed in §2.3.1–§2.3.4 are static merging methods. This means that the merged model remains the same for all samples or tasks. Given that there are differences between input samples/tasks, the model’s ability may vary when processing different samples/tasks. As shown in Figure 5 (d), some works propose to dynamically merge models (or subsets of layers) based on the samples/tasks (Li et al., 2024e; Muqeeth et al., 2024a; Tang et al., 2024b; Lu et al., 2024a; Kang et al., 2024; Ye et al., 2025) during the inference phase. EMR-Merging (Huang et al., 2024c) proposes maintaining a shared model among multiple tasks alongside a sparse task-specific model. During inference, the sample-specific sparse task model is merged with the shared dense model. In addition, for a given input, SMEAR (Muqeeth et al., 2024a) first computes a weighted average of the parameters of each expert by using the distribution of router inputs to the expert modules. The advantage of this approach is that it has a similar computational cost to that of a single expert. Twin-Merging (Lu et al., 2024a) also adaptively combines task-shared and task-private knowledge based on routing during the inference phase. Similarly, WEMoE (Tang et al., 2024b) proposes a dynamic merging Transformer architecture. It uses a standard weighted average to merge all modules except the linear layer. The linear layer is dynamically weighted and merged according to the routing network during inference. PWE MoE (Tang et al., 2024a) further extends WEMoE to a multi-objective optimization setting and uses the preference vector as input for routing. MoW-Merging (Ye et al., 2025) employs a gating network to adaptively generate merging coefficients conditioned on the input sample, thereby enabling sample-level dynamic merging and automatic classifier selection.
2.3.6. Post-calibration-based Methods.
@@
Yang et al. (2024a) introduce a post-merging method to calibrate merged models. They observed that merged models suffer from representation bias, meaning the representations extracted by the independent and merged models are very different, leading to performance degradation in the merged model. To alleviate this problem, they propose a module called ‘representation surgery’ to calibrate the representation bias. Building on this, SurgeryV2 (Yang et al., 2024b) assumes that representation bias arises at every layer of the network and therefore applies the representation surgery strategy layer-wise. ProbSurgery (Wei et al., 2025b) argues that the deterministic modeling in Surgery is suboptimal and introduces a probabilistic formulation to better eliminate representation bias. Similarly, LOT-Merging (Sun et al., 2025) adopts a layer-wise scheme that explicitly minimizes the feature drift between task-specific expert models and the merged model.
2.4. Summary and Discussion
Table 2 summarizes and compares several representative model merging methods. We make the following key observations: (1) Compared with maintaining a set of independent expert models, model merging can significantly reduce the total number of parameters that need to be stored and managed, especially when the number of tasks $N$ is large. (2) Training-free methods and training-based methods each have their own advantages. Training-free approaches are simpler and easier to deploy across different application scenarios, whereas training-based methods require additional computational resources and come with higher implementation complexity. (3) Most existing works do not require extra data at the merging stage, but the lack of data-driven merging may limit performance. In contrast, data-dependent approaches rely on labeled training or validation data, or on unlabeled test data, to guide the merging process. (4) The vast majority of model merging methods adopt a static scheme, which introduces no additional overhead at inference time. Dynamic merging, however, must first compute sample- or task-dependent merged parameters before prediction, incurring extra latency. Moreover, dynamic approaches often require storing additional expert parameters, thereby increasing memory usage. (5) The weighting granularity in model merging ranges from coarse to fine, including global, task-wise, layer-wise, and parameter-wise weighting. Finer-grained weighting schemes generally deliver better performance but typically rely on extra data to estimate appropriate merging coefficients. By contrast, many optimization-based and low-rank subspace methods are designed to work in a largely data-free manner.
Overall, there is no single “perfect” merging method that fits all scenarios; the most appropriate approach should be chosen based on performance requirements, implementation complexity, and the availability of data and computational resources.
Table 2. A brief overview of representative model merging methods. Category: The methodological category each approach belongs to. Parameters: The number of parameters in the merged model, where $1\times$ denotes the parameter size of a single-task model and NN denotes the total number of tasks. Tuning-Free: Whether the merging process is tuning-free. Required Data: What additional data (if any) are required for merging. Pattern: Whether the merging pattern is static or dynamic. Weighted Level: The granularity at which multiple models are weighted during merging. Here, =yes and =no.
Method Category Parameters Tuning-Free Required Data Pattern Weighted Level
[…]
3. Theoretical and Experimental Analysis of Model Merging
3.1. Theoretical Analysis